# Spectral Flow for the Riemann zeros

### v2

@@ VI.1 Generalized Riemann Hypothesis
Let $L(s,X)$ denote the $L$-function based on a primitive Dirichlet character $X$ of modulus $q$. Due to the complete multiplicity of the characters $X(n)X(m)=X(nm)$, it has both a series and a Euler product representation:
@@
The equations in FrancaLeClair for individual zeros on the line depend on various details such as the Gauss sum and other number-theoretic quantities, but for our purposes these are negligible for large enough $t$.gggSee for instance equation (43) in FrancaLeClair . Repeating the arguments of Section III we can propose the following.
@@
The Generalize Riemann Hypothesis is true if for $t$ large enough
@@ VI.2 Grand Riemann Hypothesis
The Grand Riemann Hypothesis is true if for $t$ large enough
@@
VII Conclusions
@@
Based on this article, the validity of the RH comes down Proposition 1, namely that the hamiltonian for the LM model and its generalizations has real eigenvalues $E_{n}(\sigma)$. We argued that this is valid due to the unitarity of the S-matrix and the reality and invertibility of the dispersion relation. In fact, it is clear from (4) that the free hamiltonian for our dispersion relation is a hermitian operator. We thus cannot see any obstructions to 1 at present. We point out that the recent work Negro seems relevant to the study of the free non-interacting part of the hamiltonian if the dispersion relation is based on the exact Riemann-Siegel $\vartheta$ function.
@@
There are many interesting avenues for further exploration. The $\{E_{n}(\sigma)\}$ are well defined for all $\sigma>\tfrac{1}{2}$, and it would be interesting to understand if they still satisfy GUE random matrix statistics as is the case for $\sigma=\tfrac{1}{2}$. We suspect that they do, since the S-matrix that defines the model is based on the primes which are pseudo-random. For the $L$-functions that violate the RH, such as the example in Section IV, it would be interesting to study if the complex eigen-energies satisfy the statistics of random non-hermitian hamiltonians classified in BL38 .

### v3

@@ VI.1 Generalized Riemann Hypothesis
Let $L(s,X)$ denote the $L$-function based on a primitive Dirichlet character $X$ of modulus $q$. The characters $X$ are periodic, $X(n)=X(n+q)$, and are all phases which are roots of unity, $X(n)=e^{i\varphi(n)}$. Thus these $L$-functions are already incorporated in the general LM model (8). Due to the complete multiplicity of the characters $X(n)X(m)=X(nm)$, it has both a series and a Euler product representation:
@@
The equations in FrancaLeClair for individual zeros on the line depend on various details such as the Gauss sum and other number-theoretic quantities, but for our purposes these are negligible for large enough $t$.iiiSee for instance equation (43) in FrancaLeClair . Repeating the arguments of Section III we can propose the following.
@@
The Generalized Riemann Hypothesis is true if for $t>t_{\star}$ large enough
@@
Here $t_{\star}$ is typically smaller than for $\zeta$.
@@ VI.2 Grand Riemann Hypothesis
The Grand Riemann Hypothesis is true if for $t$ large enough, $t>t_{\star}\approx 25$,
@@
The simplest example is the modular form based based on the Dedekind $\eta$ function:
$\eta(\tau)=q^{1/24}\prod_{n=1}^{\infty}(1-q^{n}).$ (54)
[…]
Figure 14: A plot of $-\Re(\Upsilon(s))$ for the weight $k=12$ modular $L_{f}$ based on Ramanujan $\tau$ verses $\log(t/2\pi)$ as a function of $t$ for $\sigma$ just above the critical line, $\sigma=6.01$.
VII Statistics for eigenvalues $E_{n}(\sigma)$ off the critical line.
The Montgomery-Odlyzko conjecture is that the properly normalized pair correlation function satisfies GUE statistics. In this section we compare the statistics of the actual zeros $t_{n}$ with that of $E_{n}(\sigma)$ for $\sigma>\tfrac{1}{2}$. We carry out this comparison mainly out of curiosity, and don’t present any rigorous results as far as what this entails.
Consider the normalized spacings
$\delta E_{n}^{(k)}(\sigma)=\dfrac{1}{2\pi}\left(E_{n+k}(\sigma)-E_{n}(\sigma)\right)\log\left(E_{n}(\sigma)/2\pi\right),~{}~{}~{}k=1,2,\ldots$ (55)
where $k=1$ corresponds to nearest neighbors, etc. The original conjecture concerns all pairs $k\geq 1$ and has the well-known distribution:
$1-\dfrac{\sin^{2}(\pi v)}{\pi^{2}v^{2}}=\sum_{k=1}^{\infty}\rho_{k}(v),$ (56)
where the function $\rho_{1}(v)$ gives the nearest neighbor correlation and so forth. The distribution $\rho_{1}$ is known as the Gaudin distribution, which can be expressed in terms of Fredholm determinants, however has no simple expression. It is known that a very good approximation to $\rho_{1}(v)$ is the Wigner surmise:
$\rho_{1}(v)\approx\rho_{\rm w}(v)=\dfrac{32\,v^{2}}{\pi^{2}}\,e^{-4v^{2}/\pi}.$ (57)
The above just equals the Maxwell-Boltzmann distribution of speeds $v$ in a free gas of non-relativistic particles of mass $m$, with kinetic energy $\tfrac{1}{2}mv^{2}$, at a temperature $k_{B}T=m\pi/8$.
Since we only wish to compare the statistics of the true zeros $t_{n}$ with those of $E_{n}(\sigma)$ for $\sigma>\tfrac{1}{2}$, for simplicity we only consider the nearest neighbor pair correlation $\rho_{1}(v)$ and compare it with the Wigner surmise since this requires less computational effort. In Figure 15 we present a normalized histogram for $\rho_{1}$ for $10^{5}$ eigenvalues $E_{n}(\sigma)$ for $\sigma=\tfrac{1}{2}$ which are the actual zeros $t_{n}$. The results are what we expect, namely that the results are consistent with random matrix theory. In Figure 16 we present analogous results for $\sigma=0.6$. One can readily conclude that the $E_{n}(\sigma)$ are not consistent with GUE statistics except for $\sigma=\tfrac{1}{2}$. In particular, above $\sigma=\tfrac{1}{2}$ there appears to be stronger level-repulsion since the probability density is zero in a finite range above 00 in comparison with the Wigner surmise.
Figure 15: Probability distribution for normalized consecutive eigenvalues $E_{n}(\sigma)$ for $\sigma=\tfrac{1}{2}$ in the range $10^{6}<n<10^{6}+10^{5}$. The solid curve is the Wigner surmise.
Figure 16: Probability distribution for normalized consecutive eigenvalues $E_{n}(\sigma)$ for $\sigma=0.6$ in the range $10^{6}<n<10^{6}+10^{5}$. The solid curve is the Wigner surmise.
VIII Conclusions
@@
Based on this article, the validity of the RH comes down Proposition 1, namely that the hamiltonian for the LM model and its generalizations has real eigenvalues $E_{n}(\sigma)$. We argued that this is valid due to the unitarity of the S-matrix and the reality and invertibility of the dispersion relation. In fact, it is clear from (4) that the free hamiltonian for our dispersion relation is a hermitian operator. We thus cannot see any obstructions to Proposition 1 at present. This work shows that the Hilbert-Pólya conjecture can be true if one plucks a Hamiltonian of the right feather, in particular one that leads to a scattering problem rather than a bound state problem. We point out that the recent work Negro seems relevant to the study of the free non-interacting part of the hamiltonian if the dispersion relation is based on the exact Riemann-Siegel $\vartheta$ function.
@@
There are many interesting avenues for further exploration. For the $L$-functions that violate the RH, such as the example in Section IV, it would be interesting to study if the complex eigen-energies satisfy the statistics of random non-hermitian hamiltonians classified in BL38 .